OpenStudy (itsonlycdeee):

Prove that the following identity is true: sec^4 theta - tan^4 theta = 1 + sin^2 theta / cos^2 theta

OpenStudy (souvik):

use this...\[a ^{2}-b ^{2}=(a+b)(a-b)\]

OpenStudy (souvik):

no no ....\[\sec ^{4}\theta-\tan ^{4}\theta \] =\[(\sec ^{2}\theta+\tan ^{2}\theta)(\sec ^{2}\theta-\tan ^{2}\theta)\]

OpenStudy (souvik):

you would start from left hand side...........